# Neurosymbolic Association Rule Mining from Tabular Data

### v2

@@ 2 Related Work
Deep Learning (DL) in ARM. To the best of our knowledge, very few DL-based methods can directly mine association rules from tabular data, despite DL’s widespread success. Patel and Yadav (2022) used an autoencoder (Bank et al., 2023) to mine frequent itemsets from a grocery dataset and derive association rules, but their study lacks an explicit algorithm or source code. Berteloot et al. (2024) introduced ARM-AE, another autoencoder-based ARM method. ARM-AE was not extensively evaluated and yields low-confidence rules as reported in their paper (e.g., ${\%}3$ confidence in one dataset) and our findings (Section 4.1). Karabulut et al. (2024) proposed a DL-based ARM leveraging autoencoders, however, it is tailored to Internet of Things (IoT) domain, incorporating sensor data and knowledge graphs. Note that the term rule learning encompasses different tasks, such as learning rules over graphs (Ho et al., 2018), which are out of scope for this paper.
@@ 3.2 Autoencoder Architecture and Training Stage
Let $F=\{f_{1},f_{2},...,f_{k}\}$ be a set of $k$ features in a tabular dataset (e.g. columns in a table) and $f_{i}^{1...c_{i}}$ represent categories (possible values) for feature $f_{i}$ $(1\leq i\leq k)$ where $c_{i}=|f_{i}|$ indicates the number of categories for feature $f_{i}$. To represent such a tabular dataset as an ARM problem as defined in Section 2, we define the full set of items $I$ as consisting of all possible categories of all features $I=\{f_{1}^{1},...,f_{1}^{c_{1}},...,f_{k}^{1},...,f_{k}^{c_{k}}\}$. Each transaction $T$ in the tabular dataset (e.g., a row in a table) is represented as an itemset $T\subset I$, where each feature contributes exactly one item: $\forall i\in\{1,...,k\},\exists!j\in\{1,...,c_{i}\}\text{ such that }f_{i}^{j}\in T$. Using one-hot encoded representation, $f_{i}^{1...c_{i}}=\{0,1\}$, with 0 and 1 indicating the absence or the presence of a feature category in a given transaction.
@@ 3.3 Rule Extraction Stage
Intuition. Autoencoders can learn a neural representation of data and this representation includes the associations between the feature categories. We hypothesize that the reconstruction ability of autoencoders can be used to extract these associations. After training, if a forward run on the trained model with a set of marked categories $A$ results in successful reconstruction (high probability) of categories $C$, we say that marked features $A$ imply the successfully reconstructed features $C$, such that $A\rightarrow{C}\setminus A$ (no self-implication).
@@
Input: Trained autoencoder: $AE$, max antecedents: $a$, similarity thresholds $\tau_{a},\tau_{c}$
@@
1 $\mathcal{R}\leftarrow\emptyset$, $\mathcal{F}\leftarrow$ $AE.input\_feature\_categories$;
@@ 8                   $\mathbf{p}\leftarrow AE(\mathbf{v})$;
[…]
@@
Example. Figure 2 shows rule extraction. Let $weather$ and $beverage$ be features with categories cold, warm and tea, coffee, soda, respectively. We start with a test vector of size 5, assigning equal probabilities per feature: $[0.5,0.5,0.33,0.33,0.33]$. Then we mark $weather(warm)$ by assigning 1 to $warm$ and 0 to $cold$, $[1,0,0.33,0.33,0.33]$, and call the resulting vector a test vector. Assume that after a forward run, $[0.7,0.3,0.04,0.1,0.86]$ is received as the output probabilities. Since the probability of $p_{weather(warm)}=0.7$ is bigger than the given antecedent similarity threshold ($\tau_{a}=0.5$), and $p_{beverage(soda)}=0.86$ probability is higher than the consequent similarity threshold ($\tau_{c}=0.8$), we conclude with $weather(warm)\rightarrow beverage(soda)$.
Algorithm. The rule extraction process is given in Algorithm 1. Line 1 stores input feature categories into $\mathcal{F}$. Lines 2-14 iterate over the number of antecedents $i$ and line 3 generates an $i$-feature combination $\mathcal{F}$. Lines 4-13 iterate over the feature combinations $\mathcal{C}$. For each combination $S$, line 5 creates a vector with uniform probabilities per feature category, e.g., $[0.5,0.5,0.33,0.33,0.33]$ vector in the example given above. Line 6 creates a set of test vectors where a combination of feature categories in $S$ are marked per test vector. This corresponds to the $[1,0,0.33,0.33,0.33]$ vector in the example where $weather(warm)$ was marked. Lines 7-13 iterate over the test vectors. Line 8 performs a forward run on the trained autoencoder $AE$ with each test vector $v$. Lines 9-11 compare the output probabilities corresponding to $S$ with a given antecedent similarity threshold $\tau_{a}$, and the algorithm continues with high probability $S$ values. Lines 12-13 compare the output probabilities for the categories in $\mathcal{F}$ that are not in $S$ already ($F\setminus S$), with a given consequent similarity threshold $\tau_{c}$ and stores categories with high probability. Finally, line 14 removes the low support categories from $\mathcal{F}$ so they are ignored in line 3 of the next iteration.
@@ 4.1 Experimental Setting 1: Execution Time and Rule Quality Evaluation
FP-Growth Exhaustive antecedents = 2, min_support=$\frac{1}{2}\mathbb{E}[\text{support}(R)]$, min_conf=0.8
@@ 4.2 Experimental Setting 2: Aerial+ on Downstream Tasks
Exhaustive |||| Aerial+ Exhaustive |||| Aerial+ Exhaustive |||| Aerial+
Congressional Voting Records CBA 3437 |||| 1495 91.91 |||| 92.66 0.34 |||| 0.14
BRL 2547 |||| 57 96.97 |||| 96.97 15.37 |||| 9.69
CORELS 4553 |||| 61 96.97 |||| 96.97 3.04 |||| 0.17
Mushroom CBA 27800 |||| 2785 99.82 |||| 99.82 1.75 |||| 1.30
BRL 5093 |||| 493 99.87 |||| 99.82 244 |||| 167
CORELS 23271 |||| 335 90.14 |||| 99.04 61 |||| 2
Breast Cancer CBA 695 |||| 601 66.42 |||| 71.13 0.08 |||| 0.28
BRL 2047 |||| 290 71.13 |||| 71.46 16.82 |||| 14.5
CORELS 2047 |||| 369 73.69 |||| 75.82 1.42 |||| 0.40
Chess CBA 49775 |||| 34490 94.02 |||| 93.86 24.31 |||| 6.24
BRL 19312 |||| 1518 96.21 |||| 95.93 321 |||| 119
CORELS 37104 |||| 837 81.1 |||| 93.71 106 |||| 3.87
Spambase CBA 125223 |||| 33418 84.5 |||| 85.42 23.87 |||| 7.56
BRL 37626 |||| 5190 72.78 |||| 84.93 1169 |||| 431
CORELS 275003 |||| 1409 85.37 |||| 87.28 1258 |||| 5.23
@@ Appendix A Runtime Complexity Analysis of Aerial+
    Performs a forward pass on the trained autoencoder $AE$ in $O(|\mathcal{F}|)$.
@@ Appendix C Variations of Aerial+
Input: Trained autoencoder: $AE$, max antecedents: $a$, itemset similarity $\tau_{i}$
@@
$\mathcal{I}\leftarrow\emptyset$, $\mathcal{F}\leftarrow$ $AE.input\_feature\_categories$;
@@ 12                         $\mathcal{I}\leftarrow\mathcal{I}\cup S$;
[…]
@@
Input: Trained autoencoder: $AE$, max antecedents: $a$, similarity thresholds $\tau_{a},\tau_{c}$, items of interest $\mathcal{I}_{a}$ and $\mathcal{I}_{c}$
@@
$\mathcal{R}\leftarrow\emptyset$, $\mathcal{F}\leftarrow$ $AE.input\_feature\_categories$;
@@ 17                               $\mathcal{R}\leftarrow\mathcal{R}\cup\{(S\rightarrow f)\}$;
[…]
@@ Appendix D Execution Time Experiments for Optimization-based ARM
The execution time and quality of the rules mined by the optimization-based ARM methods depend on their preset $max\_evaluation$ parameter as presented in Section 4.1. $max\_evaluation$ refers to the maximum number of fitness function evaluations, which are typically a function of rule quality metrics, before termination.
We run the optimization-based methods with the parameters described in Table 2, and with varying numbers of $max\_evaluations$. Table 3 in Section 4.1 presented the results for the Mushroom dataset. Table 6 presents the results for the remaining four datasets.
The results show that on average as the $max\_evaluations$ increases, the number of rules, execution time, and the average confidence of the rules increase while the improvement in the confidence levels stagnates after 50,000 evaluations. Optimization-based methods could not find any rules on the Spambase dataset, except the BAT algorithm, up to 100,000 evaluations which took 2̃0 minutes to terminate. The results are consistent with Mushroom dataset results given in Table 3.
@@ F.3 Training stage
$\bar{x}_{\mathcal{C}_{i}}^{(c)}\approx\mathbb{P}\left(x_{\mathcal{C}_{i}}^{(c)}=1\mid\tilde{x}\right),$ (4)
@@ F.4 Rule Extraction Stage
Forward pass and output interpretation. Next, we feed the input $A$ through the trained autoencoder to obtain output:
@@
[…]
@@ F.5 Interpretation of the formal justification
  • Probing with test vectors. Rule extraction relies on probing the decoder with a constructed input $A$ that sets $x_{i}=c$ deterministically and others to uniform noise. The resulting output $\bar{x}$ estimates $\mathbb{P}(x_{j}=c^{\prime}\mid x_{i}=c)$ (Equation 6), enabling the discovery of association rules.

### v3

@@ 2 Related Work
[…]
@@ 3.2 Autoencoder Architecture and Training Stage
Let $F=\{f_{1},f_{2},...,f_{k}\}$ be a set of $k$ features in a tabular dataset (e.g. columns in a table) and $f_{i}^{1...c_{i}}$ represent categories (possible values) for feature $f_{i}$ $(1\leq i\leq k)$ where $c_{i}=|f_{i}|$ indicates the number of categories for feature $f_{i}$. To represent such a tabular dataset as an ARM problem as defined in Section 2, we define the full set of items II as consisting of all possible categories of all features $I=\{f_{1}^{1},...,f_{1}^{c_{1}},...,f_{k}^{1},...,f_{k}^{c_{k}}\}$. Each transaction $T$ in the tabular dataset (e.g., a row in a table) is represented as an itemset $T\subset I$, where each feature contributes exactly one item: $\forall i\in\{1,...,k\},\exists!j\in\{1,...,c_{i}\}\text{ such that }f_{i}^{j}\in T$. Using one-hot encoded representation, $f_{i}^{1...c_{i}}=\{0,1\}$, with 0 and 1 indicating the absence or the presence of a feature category in a given transaction.
@@ 3.3 Rule Extraction Stage
Intuition. Autoencoders can learn a neural representation of data and this representation includes the associations between the feature categories. We hypothesize that the reconstruction ability of autoencoders can be used to extract these associations. After training, if a forward run on the trained model with a set of marked categories AA results in successful reconstruction (high probability) of categories $C$, we say that marked features AA imply the successfully reconstructed features $C$, such that $A\rightarrow{C}\setminus A$ (no self-implication).
@@
Input: Trained autoencoder: AEAE, max antecedents: $a$, similarity thresholds $\tau_{a},\tau_{c}$
@@
1 $\mathcal{R}\leftarrow\emptyset$, $\mathcal{F}\leftarrow$ AE.input_feature_categoriesAE.input\_feature\_categories;
@@ 8       $\mathbf{p}\leftarrow AE(\mathbf{v})$;
16 $\mathcal{F}\leftarrow\{f\in\mathcal{F}\mid f.\text{low\_support}=\textbf{False}\}$;
@@
Example. Figure 2 shows rule extraction. Let weatherweather and beveragebeverage be features with categories cold, warm and tea, coffee, soda, respectively. We start with a test vector of size 5, assigning equal probabilities per feature: $[0.5,0.5,0.33,0.33,0.33]$. Then we mark $weather(warm)$ by assigning 1 to warmwarm and 0 to coldcold, $[1,0,0.33,0.33,0.33]$, and call the resulting vector a test vector. Assume that after a forward run, $[0.7,0.3,0.04,0.1,0.86]$ is received as the output probabilities. Since the probability of $p_{weather(warm)}=0.7$ is bigger than the given antecedent similarity threshold ($\tau_{a}=0.5$), and $p_{beverage(soda)}=0.86$ probability is higher than the consequent similarity threshold ($\tau_{c}=0.8$), we conclude with $weather(warm)\rightarrow beverage(soda)$.
Algorithm. The rule extraction process is given in Algorithm 1. Line 1 stores input feature categories into $\mathcal{F}$. Lines 2-14 iterate over the number of antecedents $i$ and line 3 generates an $i$-feature combination $\mathcal{F}$. Lines 4-13 iterate over the feature combinations $\mathcal{C}$. For each combination $S$, line 5 creates a vector with uniform probabilities per feature category, e.g., $[0.5,0.5,0.33,0.33,0.33]$ vector in the example given above. Line 6 creates a set of test vectors where a combination of feature categories in $S$ are marked per test vector. This corresponds to the $[1,0,0.33,0.33,0.33]$ vector in the example where $weather(warm)$ was marked. Lines 7-13 iterate over the test vectors. Line 8 performs a forward run on the trained autoencoder AEAE with each test vector $v$. Lines 9-11 compare the output probabilities corresponding to $S$ with a given antecedent similarity threshold $\tau_{a}$, and the algorithm continues with high probability $S$ values. Lines 12-13 compare the output probabilities for the categories in $\mathcal{F}$ that are not in $S$ already ($F\setminus S$), with a given consequent similarity threshold $\tau_{c}$ and stores categories with high probability. Finally, line 14 removes the low support categories from $\mathcal{F}$ so they are ignored in line 3 of the next iteration.
@@ 4.1 Experimental Setting 1: Execution Time and Rule Quality Evaluation
[…]
@@ 4.2 Experimental Setting 2: Aerial+ on Downstream Tasks
Exhaustive $|$ Aerial+ Exhaustive $|$ Aerial+ Exhaustive $|$ Aerial+
Congressional Voting Records CBA 3437 $|$ 1495 91.91 $|$ 92.66 0.34 $|$ 0.14
BRL 2547 $|$ 57 96.97 $|$ 96.97 15.37 $|$ 9.69
CORELS 4553 $|$ 61 96.97 $|$ 96.97 3.04 $|$ 0.17
Mushroom CBA 27800 $|$ 2785 99.82 $|$ 99.82 1.75 $|$ 1.30
BRL 5093 $|$ 493 99.87 $|$ 99.82 244 $|$ 167
CORELS 23271 $|$ 335 90.14 $|$ 99.04 61 $|$ 2
Breast Cancer CBA 695 $|$ 601 66.42 $|$ 71.13 0.08 $|$ 0.28
BRL 2047 $|$ 290 71.13 $|$ 71.46 16.82 $|$ 14.5
CORELS 2047 $|$ 369 73.69 $|$ 75.82 1.42 $|$ 0.40
Chess CBA 49775 $|$ 34490 94.02 $|$ 93.86 24.31 $|$ 6.24
BRL 19312 $|$ 1518 96.21 $|$ 95.93 321 $|$ 119
CORELS 37104 $|$ 837 81.1 $|$ 93.71 106 $|$ 3.87
Spambase CBA 125223 $|$ 33418 84.5 $|$ 85.42 23.87 $|$ 7.56
BRL 37626 $|$ 5190 72.78 $|$ 84.93 1169 $|$ 431
CORELS 275003 $|$ 1409 85.37 $|$ 87.28 1258 $|$ 5.23
@@ Appendix A Runtime Complexity Analysis of Aerial+
    Performs a forward pass on the trained autoencoder AEAE in $O(|\mathcal{F}|)$.
@@ Appendix C Variations of Aerial+
Input: Trained autoencoder: AEAE, max antecedents: $a$, itemset similarity $\tau_{i}$
@@
[…]
@@ 12          $\mathcal{I}\leftarrow\mathcal{I}\cup S$;
14 $\mathcal{F}\leftarrow\{f\in\mathcal{F}\mid f.\text{low\_support}=\textbf{False}\}$;
@@
Input: Trained autoencoder: AEAE, max antecedents: $a$, similarity thresholds $\tau_{a},\tau_{c}$, items of interest $\mathcal{I}_{a}$ and $\mathcal{I}_{c}$
@@
[…]
@@ 17             $\mathcal{R}\leftarrow\mathcal{R}\cup\{(S\rightarrow f)\}$;
19 $\mathcal{I}_{a}\leftarrow\{f\in\mathcal{I}_{a}\mid f.\text{low\_support}=\textbf{False}\}$;
@@ Appendix D Execution Time Experiments for Optimization-based ARM
The execution time and quality of the rules mined by the optimization-based ARM methods depend on their preset max_evaluationmax\_evaluation parameter as presented in Section 4.1. max_evaluationmax\_evaluation refers to the maximum number of fitness function evaluations, which are typically a function of rule quality metrics, before termination.
We run the optimization-based methods with the parameters described in Table 2, and with varying numbers of max_evaluationsmax\_evaluations. Table 3 in Section 4.1 presented the results for the Mushroom dataset. Table 6 presents the results for the remaining four datasets.
The results show that on average as the max_evaluationsmax\_evaluations increases, the number of rules, execution time, and the average confidence of the rules increase while the improvement in the confidence levels stagnates after 50,000 evaluations. Optimization-based methods could not find any rules on the Spambase dataset, except the BAT algorithm, up to 100,000 evaluations which took 2̃0 minutes to terminate. The results are consistent with Mushroom dataset results given in Table 3.
@@ F.3 Training stage
[…]
@@ F.4 Rule Extraction Stage
Forward pass and output interpretation. Next, we feed the input AA through the trained autoencoder to obtain output:
@@
$\bar{x}_{\mathcal{C}_{j}}^{(c^{\prime})}\geq\mathcal{T}_{c},\quad\text{for }j\neq i\text{ (no self-implication)},$
@@ F.5 Interpretation of the formal justification
  • Probing with test vectors. Rule extraction relies on probing the decoder with a constructed input AA that sets $x_{i}=c$ deterministically and others to uniform noise. The resulting output $\bar{x}$ estimates $\mathbb{P}(x_{j}=c^{\prime}\mid x_{i}=c)$ (Equation 6), enabling the discovery of association rules.